OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10 -2y + 10 2y + 10 2y - 10 -2y - 10

jimthompson5910 (jim_thompson5910):

they want you to solve -x - 2y = -10 for x

OpenStudy (anonymous):

what do I plug in for it?

jimthompson5910 (jim_thompson5910):

nothing, you're just isolating x in that equation

jimthompson5910 (jim_thompson5910):

example if you had x+y = 10 isolating x would give you x+y = 10 x+y-y = 10-y x+0y = 10-y x = 10-y

OpenStudy (anonymous):

so 2y-10?

jimthompson5910 (jim_thompson5910):

-x - 2y = -10 -x - 2y+2y = -10+2y .. Add 2y to both sides. -x + 0y = -10 + 2y -x = -10 + 2y what's next?

OpenStudy (anonymous):

divide -1 10-2y

jimthompson5910 (jim_thompson5910):

good x = 10 - 2y x = 10 + (-2y) x = -2y + 10

OpenStudy (anonymous):

okay thank you!

jimthompson5910 (jim_thompson5910):

np